# Subsystem localization

### v2

@@ V.1 Toy model
Figure 12: A localization measure $C^{A}$ vs. $V$ plots for the mid-spectrum state of $H_{T}$ for different $t^{\prime}$. One inset shows the probability of a particle being at subsystem $p_{A}$ vs. $V$ for the same state. The top-right inset shows the scaling of $V^{*}$ (the $V$ for which $C^{A}$ attains a peak) with $t^{\prime}$.
Here, we propose a minimal toy model to help understand some of the numerical results obtained for the Hamiltonian in Eq. (1). One surprising observation from our numerical simulations is the reentrance of the delocalized phase with increasing $V$. Typically, at least for $t^{\prime}=0$, increasing $V$ tends to promote localization in subsystem B. It is therefore natural to expect that introducing a nonzero $t^{\prime}$ would further localize subsystem A as $V$ increases. Contrary to this expectation, we find that beyond a certain threshold (that depends on $t^{\prime}$), increasing $V$ leads to the reappearance of the delocalized phase. To investigate this phenomenon, we study a minimal toy model described by the Hamiltonian (using the same site indexing as shown in the schematic Fig. 1),
$H_{T}=-(c^{\dagger}_{1}c_{2}+t^{\prime}c_{1}^{\dagger}c_{3}+t^{\prime}c_{2}^{\dagger}c_{4}+hc)+V(c^{\dagger}_{3}c_{3}-c^{\dagger}_{4}c_{4}),\$ (10)
with the aim of capturing the mechanism behind the reentrance of the delocalized phase. If $t^{\prime}=0$, the A and B parts of the system are completely decoupled, and the eigenstates (specifically, two of them) remain entirely on the B part of the system and are fully localized. In contrast, for the other two eigenstates that lie entirely within the A part, the probability of finding the particle on sites $1$ and $2$ remains equal. If one introduces a finite $t^{\prime}$ and investigates how this scenario evolves with varying $V$, to quantify the degree of delocalization of an eigenstate within subsystem A (which consists only of sites $1$ and $2$), we define the following measure for an eigenstate $|\xi\rangle$:
@@
Now, among the two sites in A, if $|\xi\rangle$ is completely localized on only one of the sites (either at site $1$ or $2$), then $C^{A}=1$, and on the other hand, if it is completely delocalized (probabilities to be at site $1$ and $2$ are equal), then $C^{A}=0$. Given $H_{T}$ is a $4\times 4$ matrix, $C^{A}$ can be obtained analytically, and for the mid-spectrum state,
@@
Figure 12 (main panel) shows the variation of $C^{A}$ with $V$ for different values of $t^{\prime}$. As $V$ increases, $C^{A}$ initially increases, reaches a peak at $V=V^{*}$, and then decreases as $V$ continues to grow. This behavior closely mimics the reentrance phenomenon of delocalization observed in the original Hamiltonian $H$. We also obtain the analytical expression $V^{*}=\frac{t^{\prime}}{\sqrt{2}}$ by solving the equation $\frac{dC^{A}}{dV}\big{|}_{V^{*}}=0$.
@@
as a function of $V$. We find that as $V$ increases, this probability approaches $1$, indicating that the state becomes increasingly localized within subsystem A, making it a significant state for our analysis of subsystem A.
@@ V.2 Effective model for the subsystem
We will see here that a model closely related to ours (Eq. 1) can be reduced to a simple effective model for the subsystem. The model we consider here does not have the hopping term along the $B$ leg (the bath). The Hamiltonian of this model is the same as the one appears in Eq. 1 with $t_{B}=0$. The two models are expected to give similar results in the large $t^{\prime}$ and $V$ limits.
Consider that $\psi=\sum_{j=1}^{N}a_{j}\ket{A_{j}}+\sum_{i^{\prime}=1}^{N}b_{i^{\prime}}\ket{B_{i^{\prime}}}$ is an eigenstate of the Hamiltonian of the current model (where $t_{B}=0$). Here |$\ket{A_{j}}$ and |$\ket{B_{i^{\prime}}}$ are the site basis of subsystem $A$ and bath $B$, respectively, and $a_{j}$ and $b_{i^{\prime}}$ are the corresponding coefficients. It may be noted that the index $i^{\prime}$ corresponds to $N+i$ in the original Hamiltonian in Eq. 1. If $E$ is the eigevalue associated with $\psi$, then
$-t^{\prime}a_{n}+V_{n}b_{n}=$ $Eb_{n}$ (14)
$-t_{A}a_{n+1}-t_{A}a_{n-1}-t^{\prime}b_{n}=$ $Ea_{n}$ (15)
Here $V_{n}=V\cos{(2\pi\beta n+\phi)}$, the AA potential as appears in Eq. 1. From Eq. 14, we get:
$b_{n}=\frac{-t^{\prime}a_{n}}{(E-V_{n})}.$
We use this expression to replace $b_{n}$ in Eq. 15; this yields:
$-t_{A}a_{n+1}-t_{A}a_{n-1}+\frac{(t^{\prime})^{2}a_{n}}{(E-V_{n})}=Ea_{n}.$ (16)
@@
$V^{eff}_{n}=\frac{\lambda\cos{(2\pi\beta n+\phi)}}{1-\lambda\alpha\cos{(2\pi\beta n+\phi)}},$
@@ VI Conclusion
In summary, in this work, we study a two-leg ladder model, where one leg (called bath) is described by an AA-type Hamiltonian (which can undergo a localization-delocalization transition by tuning the strength of the incommensurate potential) and the other leg by the usual nearest-neighbor tight-binding model, which we refer to as the subsystem. By introducing the coupling $t^{\prime}$ between these two legs, we obtain a rich phase diagram as shown schematically in Fig. 2. We notice that beyond a critical coupling strength $t^{\prime}_{c}$, there exists a regime where the subsystem can be completely localized. Also, we find different parameter regimes where the transport can be ballistic, subdiffusive, and superdiffusive in nature. However, it is important to note that there is an intermediate region, labeled as P4 phase in Fig. 2, where the subsystem exhibits a complex phase structure. Here it may show super- or subdiffusive behavior based on parameter values and time scale of investigation. A more detailed study is required to gain better insight into the nature of phases in this crossover region. We further explain some of these results using a toy model and also by mapping it to a well-studied 1D GAA model.
@@ Appendix A Results for random cosine model
Instead of the correlated AA type disorder, if we take a random disorder in bath $B$, then the phase structure of the subsystem $A$ is expected to be completely different. To verify this, we replace the uniform phase factor $\phi$ in Eq. 1 by a random site-dependent phase $\phi_{i}$. We then calculate $\langle PR_{A}\rangle$ (averaged over multiple realizations of $\phi_{i}$) as function of the subsystem size ($N$). The results may be found in Fig. 13. Fitting of the data by $\langle PR_{A}\rangle\sim N^{\eta}$ shows that the exponent ($\eta$) is very small regardless of the value of $V$. This indicates that a random (uncorrelated) potential in bath $B$ localizes all states in the subsystem $A$; this is in accordance with what is expected from the Anderson localization in one-dimentional system [1].
Appendix B Scaling of projected probability distribution
@@
Appendix C Understanding Phases for $t^{\prime}<t^{\prime}_{c}$
We note that our subsystem $A$ is in a fully delocalized (or in ballistic) phase when $t^{\prime}=0$ (irrespective of the value of $V$). We also have analyzed phases of the system when $t^{\prime}>t^{\prime}_{c}$ ($t^{\prime}_{c}\approx 4.6$). In this Appendix, we present results for $0<t^{\prime}<t^{\prime}_{c}$. For definiteness, we focus on results corresponding to $t^{\prime}=3$.
Figure 16: $\langle PR_{A}\rangle$ vs. N plot for $t^{\prime}=3$ at $V$ = 1 and $V$ = 10.
We have calculated $IPR_{A}$ for $\epsilon=0.5$, and $t^{\prime}=3$; we do not observe any localized region in the intermediate V values. Corresponding results can be found in Fig. 15. A similar behavior has been observed for $t^{\prime}=4$. We have further analyzed the scaling behavior of disorder-averaged $\langle PR_{A}\rangle$ at different V regions. The $\langle PR_{A}\rangle$ vs. N plot 16 suggests that at very low V, the subsystem A is completely delocalized, but if we increase V, we will see only a few states at the edge that are becoming strongly localized. However, the middle states always remain delocalized (even in larger V). The scaling exponent of $\langle PR_{A}\rangle$ reveals that the middle states do not have the same degree of delocalization at $V=1$ and $V=10$. States at $V=10$ are much weakly delocalized compared to $V=1$.
We have also investigated the dynamical properties of the subsystem for $t^{\prime}=3$. Fig. 17 shows for $V=1$, $\gamma\sim 2$, which is a signature of ballistic growth; on the other hand, unlike $t^{\prime}=5$, in this case for $V=10$, we see a subdiffusive growth as the exponent $\gamma=0.76$. In the intermediate regime, we also find a superdiffusive regime (see $V=5$ and $V=7$ data), where $21<\gamma<$ 2.

### v3

@@ V.1 Toy model
Figure 15: A localization measure $C^{A}$ vs. $V$ plots for the most significant state of $H_{T}$ for different $t^{\prime}$. The top-left inset shows the probability of a particle being at subsystem $p_{A}$ vs. $V$ for the same state. The top-right inset shows the scaling of $V^{*}$ (the $V$ for which $C^{A}$ attains a peak) with $t^{\prime}$.
Here, we propose a minimal toy model of four sites to qualitatively understand some of the numerical results obtained for the Hamiltonian in Eq. (1). Primarily, we make three observations based on our numerical results when $t^{\prime}>t^{\prime}_{c}$: (1) the subsystem becomes localized over an intermediate range of $V$, (2) the extent of this localization regime ($V_{1}<V<V_{2}$) increases with $t^{\prime}$, and perhaps most surprisingly (3) we observe a reentrance of the delocalized phase in the subsystem at larger $V$ (as discussed in Sec. III.6).
One of our goals here is to investigate whether similar phenomena can also be observed in a minimal, fully analytically solvable toy model consisting of only four sites. In this setup, sites $1$ and $2$ constitute the subsystem $A$, while sites $3$ and $4$ form the bath $B$ (see Fig. 1 with $N=2$). Localization in $B$ is induced by assigning on-site potentials +V+V and V-V to sites $3$ and $4$, respectively. To ensure complete localization in $B$ for any finite $V$, we set the hopping amplitude between sites $3$ and $4$ to zero. In the absence of coupling between $A$ and $B$ (i.e., for $t^{\prime}=0$), this guarantees full localization within $B$.
The minimal toy model is described by the Hamiltonian,
$H_{T}=-(c^{\dagger}_{1}c_{2}+t^{\prime}c_{1}^{\dagger}c_{3}+t^{\prime}c_{2}^{\dagger}c_{4}+hc)+V(c^{\dagger}_{3}c_{3}-c^{\dagger}_{4}c_{4}).\\$ (15)
If $t^{\prime}=0$, the A and B parts of the system are completely decoupled. The two eigenstates in B are fully localized, and the two in A are maximally delocalized. For those maximally delocalized eigenstates that lie in A, the probability of finding the particle on sites $1$ and $2$ remains equal. Next, we introduce a finite $t^{\prime}$ and investigate how this scenario evolves with varying $V$. To quantify the degree of localization of an eigenstate within subsystem A, we define the following measure for an eigenstate $|\xi\rangle$:
@@
Now, among the two sites in A, if $|\xi\rangle$ is completely localized on only one of the sites (either at site $1$ or $2$), then $C^{A}=1$, and on the other hand, if it is completely delocalized (probabilities to be at site $1$ and $2$ are equal), then $C^{A}=0$. Given $H_{T}$ is a $4\times 4$ matrix, $C^{A}$ can be obtained analytically, and for the most significant state,
@@
Figure 15 (main panel) shows the variation of $C^{A}$ with $V$ for different values of $t^{\prime}$. As $V$ increases, $C^{A}$ initially increases, reaches a peak at $V=V^{*}$. We obtain the analytical expression $V^{*}=\frac{t^{\prime}}{\sqrt{2}}$ by solving the equation $\frac{dC^{A}}{dV}\big|_{V^{*}}=0$. It is also straightforward to check that $C^{A}\to 1$ as $V\to V^{*}$, indicating a completely localized phase in subsystem A. For $V>V^{*}$, however, $C^{A}$ begins to decrease. This behavior closely mirrors the emergence of a localized phase at intermediate $V$ and the subsequent reentrance of the delocalized phase in the large-$V$ regime, as observed in the original Hamiltonian $H$. Moreover, we find that the decay of $C^{A}$ for $V>V^{*}$ becomes progressively slower with increasing $t^{\prime}$, suggesting that the localized phase persists over a broader range of $V$ as $t^{\prime}$ increases. This trend is consistent with our earlier observations for the ladder Hamiltonian.
@@
as a function of $V$. We find that as $V$ increases, this probability approaches $1$, indicating that the state becomes increasingly significant within subsystem A.
Moreover, we have also studied a bond-localization measure, defined as $\Omega_{\xi}=4(|\langle 1|\xi\rangle|^{2}|\langle 3|\xi\rangle|^{2}+|\langle 2|\xi\rangle|^{2}|\langle 4|\xi\rangle|^{2}$), for our toy model. This measure approaches 11 ($0.5$) when the eigenstate is completely bond localized (delocalized). For a fixed $t^{\prime}$, we find that $\Omega_{\xi}$ approaches $1$ within an intermediate range of $V$, signaling the onset of bond localization. However, upon further increasing $V$, $\Omega_{\xi}$ decreases again. This observation reinforces the connection between re-entrant behavior and bond localization, even within our minimal toy model.
@@ V.2 Effective model for the subsystem
Here, we show that a model closely related to ours (Eq. 1) can be reduced to a simple effective one-dimensional model for the subsystem. The model we consider here does not have the hopping term along the $B$ leg (the bath). The Hamiltonian of this model is the same as the one appears in Eq. 1 with $t_{B}=0$. The two models are expected to give similar results in the large $t^{\prime}$ and $V$ limits.
[…]
$-t^{\prime}a_{i}+V_{i}b_{i}=$ $Eb_{i}$ (19)
$-t_{A}a_{i+1}-t_{A}a_{i-1}-t^{\prime}b_{i}=$ $Ea_{i}$ (20)
Here $V_{i}=V\cos{(2\pi\beta i+\phi)}$, the AA potential as appears in Eq. 1. From Eq. 19, we get:
$b_{i}=\frac{-t^{\prime}a_{i}}{(E-V_{i})}.$
We use this expression to replace $b_{i}$ in Eq. 20; this yields:
$t_{A}a_{i+1}-t_{A}a_{i-1}+\frac{(t^{\prime})^{2}a_{i}}{(E-V_{i})}=Ea_{i}.$ (21)
@@
$V^{eff}_{i}=\frac{\lambda\cos{(2\pi\beta i+\phi)}}{1-\lambda\alpha\cos{(2\pi\beta i+\phi)}},$
@@ VI Conclusion
In summary, in this work, we study a two-leg ladder model, where one leg (called bath) is described by an AA-type Hamiltonian (which can undergo a localization-delocalization transition by tuning the strength of the incommensurate potential) and the other leg by the usual nearest-neighbor tight-binding model, which we refer to as the subsystem. By introducing the coupling $t^{\prime}$ between these two legs, we obtain a rich phase diagram as shown schematically in Fig. 2. We notice that beyond a critical coupling strength $t^{\prime}_{c}$, there exists a regime where the subsystem can be completely localized. Also, we find different parameter regimes where the transport in the subsystem can be ballistic, subdiffusive, and superdiffusive in nature. However, it is important to note that there is an intermediate region, labeled as P4 phase in Fig. 2, where the subsystem exhibits an anomalous phase structure. Here it may show super- or subdiffusive behavior based on parameter values and time scale of investigation. A more detailed study is required to gain better insight into the nature of phases in this crossover region. We further explain some of these results using a toy model and also by mapping it to a well-studied 1D GAA model.
Within the scope of this work, we could not study the nature of different phase transitions, associated scaling exponents and the universality classes. One of the reasons being that we do not have precise analytical (or even numerical) estimation for the phase boundaries. Further studies are required to address these issues.
@@ Appendix A Results for random cosine model
Instead of the correlated AA type disorder, if we take a random disorder in bath $B$, then the phase structure of the subsystem $A$ is expected to be completely different. To verify this, we replace the uniform phase factor $\phi$ in Eq. 1 by a random site-dependent phase $\phi_{i}$. We then calculate $\langle PR_{A}\rangle$ (averaged over multiple realizations of $\phi_{i}$) as function of the subsystem size ($N$). The results may be found in Fig. 16. Fitting of the data by $\langle PR_{A}\rangle\sim N^{\eta}$ shows that the exponent ($\eta$) is very small regardless of the value of $V$. This indicates that a random (uncorrelated) potential in bath $B$ localizes all states in the subsystem $A$; this is in accordance with what is expected from the Anderson localization in one-dimensional system [1].
Figure 18: Log-Log plots for $\langle PR_{A}\rangle$ vs. NN with $t^{\prime}=3$. The scaling exponent $\eta$ is calculated for $V$ = 1 and $V$ = 10.
Appendix B Understanding Phases for $t^{\prime}<t^{\prime}_{c}$
In this Appendix, we present results for $t^{\prime}<t^{\prime}_{c}\approx 4.4$. For definiteness, we focus on results corresponding to $t^{\prime}=3$. The results for $t^{\prime}/V\ll 1$ (P5 phase) will be discussed separately in Appendix F.
We have calculated $IPR_{A}$ for $\epsilon=0.5$, and $t^{\prime}=3$; we do not observe any localized region in the intermediate $V$ values. Corresponding results can be found in Fig. 17. We have further analyzed the scaling behavior of $\langle PR_{A}\rangle$ in different $V$ regions. The $\langle PR_{A}\rangle$ vs. $N$ plot (Fig. 18) suggests that at very low $V$, the subsystem A is completely delocalized with $\eta\approx 1$, but if we increase $V$, we find some admixture of localized and delocalized states with $\eta<1$.
We have also investigated the dynamical properties of the subsystem for $t^{\prime}=3$. Fig. 19 shows for $V=1$, $\gamma\sim 2$, which is a signature of ballistic growth; on the other hand, in this case for $V=10$ and $15$, we see a subdiffusive growth as the exponent $\gamma=0.76$. In the intermediate regime, we also find a superdiffusive property (see $V=5$ and $V=7$ data), where $1<\gamma<2$. In this context, see our discussion in Appendix G for more details.
Appendix C Energy-resolved IPR for subsystem, bath and full system
To investigate possible correlations between the energy scales of the states and their localization properties, we plot the energy-resolved IPR for subsystem $A$, bath $B$, and the full system $A+B$. For subsystem $A$, we compute $IPR_{A}$ as defined in Eq. 5; analogously, we evaluate $IPR_{B}$ for the bath. For the full system, we calculate the IPR corresponding to its actual eigenkets (rather than their projections). The corresponding results are presented in Fig. 20.
Figure 20: (a)-(c): Plots of energy-resolved $IPR_{A}$ as function of $V$ for subsystem $A$. (d)-(f): Plots of $IPR_{B}$ as function of $V$ for bath $B$. (g)-(i): Plots of energy-resolved IPRIPR as function of $V$ for full system $A+B$. Calculations are performed for the total subsystem size $N=500$ and for $t^{\prime}$ = 1, 5 and 10. For calculations of $IPR_{A}$ and $IPR_{B}$, we take only $\epsilon=0.5$ significant states.
[…]
In analyzing the localization phase diagram of the subsystem in the thermodynamic limit, the concept of a mobility edge does not play particularly any important role. In fact, the scaling relation, $\langle PR_{A}\rangle\sim N^{\eta}$, remains valid even when the subsystem shows mobility edges. To understand this, we write $\langle PR_{A}\rangle$ as a sum of two different averages:
$\langle PR_{A}\rangle=f_{D}\langle PR_{A}\rangle_{D}+f_{L}\langle PR_{A}\rangle_{L},$ (23)
where $\langle PR_{A}\rangle_{D}$ is the $PR_{A}$ averaged over delocalized or weakly delocalized $\epsilon$-significant states, and similarly, $\langle PR_{A}\rangle_{D}$ is the $PR_{A}$ averaged over localized $\epsilon$-significant states. In the equation, $f_{D}=|S_{\epsilon,D}|/|\mathcal{S}_{\epsilon}|$ and $f_{L}=|S_{\epsilon,L}|/|\mathcal{S}_{\epsilon}|$. We recall that $\mathcal{S}_{\epsilon}$ is the set of $\epsilon$-significant states and $|\mathcal{S}_{\epsilon}|$ denotes the number of elements in the set. Similarly, $S_{\epsilon,D}$ and $S_{\epsilon,L}$ represent, respectively, the set of delocalized (including weakly delocalized) states and the set of localized states. The number of elements in them are, respectively, represented by $|S_{\epsilon,D}|$ and $|S_{\epsilon,L}|$. Since $|S_{\epsilon}|=|S_{\epsilon,D}|+|S_{\epsilon,L}|$, we have
$f_{D}+f_{L}=1,$ (24)
with $f_{D},f_{L}\geq 0$.
In the thermodynamic limit, $\langle PR_{A}\rangle_{L}=O(1)$ and $\langle PR_{A}\rangle_{D}=O(N^{\eta})$, where $0<\eta\leq 1$. Now, due to the condition in Eq. 24, we can have the following three situations in the thermodynamic limits: (a) $f_{D}=0$ and $f_{L}=1$, (b) $f_{D}=1$ and $f_{L}=0$, and (c) $0<f_{D},f_{L}<1$. In the first and second cases, the subsystem will be in the localized and delocalized (weakly delocalized) phases, respectively. The third case is an interesting case where the subsystem will have both localized and delocalized (weakly delocalized) states. This case also accounts for the situation where the subsystem has mobility edges. In the last two cases (i.e., whenever $f_{D}>0$), $\langle PR_{A}\rangle\sim N^{\eta}$, as we hypothesized at the beginning. From the log-log plot of $\langle PR_{A}\rangle$ vs $N$, one can extract the exponent $\eta$ and also estimate the fraction of delocalized (including weakly delocalized) states $f_{D}$. The value of $\eta$ helps determine if the subsystem is in delocalized, localized, or in weakly delocalized (fractal) phase.
Appendix D A measure for bond-localization
A general eigenket of the full system is given by
[…]
where $\sum_{i=1}^{N}\left(|a_{i}|^{2}+|a_{N+i}|^{2}\right)=1$. We recall that the $i$th site in the subsystem $A$ is directly connected to the $(N+i)$th site in the bath $B$ by the hopping parameter $t^{\prime}$. A straight forward calculation of the bond localization parameter (as defined in Eq. 9) shows that
$\kappa_{\xi}=4\sum_{i=1}^{N}\left[Re(a_{i}^{*}a_{N+i})\right]^{2}.$ (26)
It is not difficult to see that the value of $\kappa_{\xi}$ is bounded by 1. This can be shown in the following way.
$\begin{split}\kappa_{\xi}&\leq 4\sum_{i=1}^{N}|a_{i}^{*}a_{N+i}|^{2}\\ &=4\sum_{i=1}^{N}|a_{i}|^{2}|a_{N+i}|^{2}\\ &\leq 4(\sum_{i=1}^{N}|a_{i}|^{2})(\sum_{i=1}^{N}|a_{N+i}|^{2})\\ &\leq 4\left[\frac{1}{2}\left(\sum_{i=1}^{N}|a_{i}|^{2}+\sum_{i=1}^{N}|a_{N+i}|^{2}\right)\right]^{2}\\ &=1.\end{split}$ (27)
In the last line, we have used the normalization condition, $\sum_{i=1}^{N}\left(|a_{i}|^{2}+|a_{N+i}|^{2}\right)=1$.
We now determine the values of $\kappa_{\xi}$ for different types of states.
  1. 1.
[…]
  2. 2.
    Delocalized state in subsystem $A$: Consider a state that is fully delocalized in $A$ but does not have presence in $B$. For this state $|a_{i}|=\frac{1}{\sqrt{N}}$ and $a_{N+i}=0$ for all $i$ in Eq. 25. For this state, $\kappa_{\xi}=0$.
  3. 3.
    Delocalized state in full system: consider a state that is fully delocalized over the full system. For this state $|a_{i}|=|a_{N+i}|=\frac{1}{\sqrt{2N}}$ for any $i$ in Eq. 25. For this state, $\kappa_{\xi}=\frac{1}{N}$, which vanishes in the thermodynamic limit.
@@ Appendix D A measure for bond-localization
Figure 22: Contour plot of the scaling exponent $\gamma$ as a function of $t^{\prime}$ and $V$, where we restrict ourselves to small $t^{\prime}$ regime.
Figure 23: Show the system size dependence on the exponent $\gamma$ for different phase.
Figure 24: Show the variation of the exponent $\gamma$ with $V$ in each different phases.
Appendix E Scaling of projected probability distribution
@@
Appendix F P5 phase: A contour plot for scaling exponent of $\gamma$
In Fig. 22, we present a zoomed-in view of the small-$t^{\prime}/V\ll 1$ regime of Fig. 12 (the contour plot of $\gamma$ in the $t^{\prime}$$V$ plane). This figure clearly shows a crossover from super-diffusive to ballistic behavior in the small-$t^{\prime}$ regime. We recall that the subsystem $A$ shows ballistic behavior in $t^{\prime}\to 0$ limit. We identify this as P5 phase. The extent of this P5 phase grows with increasing $V$, suggesting that in the $V\to\infty$ limit, the system approaches the ballistic phase for any finite $t^{\prime}$. This behavior is intuitively expected: in this limit, the coupling between subsystem $A$ and bath $B$ becomes irrelevant, and the dynamics is effectively governed solely by $H_{A}$ provided we initiate the dynamics in the subsystem $A$.
Appendix G P4 phase: Strong dependence on system size and model parameters
Here, we discuss the strong system-size and model parameter dependence of the dynamical scaling exponent $\gamma$. Figure 23 shows the variation of $\sigma^{2}$ with time $t$ for different system sizes across various phases. We find that the exponent $\gamma$ remains essentially unchanged for $N=200$, $1000$, and $2000$ in the $P1$ (ballistic), $P3$ (sub-diffusive), and $P5$ (ballistic/super-diffusive) phases. However, in the $P4$ phase, the variation of $\gamma$ with system size is significant. In fact, as shown in Fig. 23, the exponent extracted for $N=200$ suggests sub-diffusive behavior, whereas for $N=1000$ and $2000$ it indicates super-diffusive behavior for the same choice of parameters within the $P4$ regime. This makes it extremely difficult to unambiguously determine the dynamical character of this phase.
Moreover, within a stable phase, one expects the system to exhibit similar behavior in the neighborhood of any given point in parameter space. This is indeed what we observe in Fig. 24 for the P1 (ballistic), P3 (sub-diffusive), and P5 (ballistic/super-diffusive) phases, where all nearby points display consistent dynamical exponents. In contrast, within the P4 phase, we find non-monotonic behavior of $\gamma$ with respect to a small change in a parameter. To illustrate this more clearly, we plot $\sigma^{2}$ as a function of $t$ for $V=9$, $10$, and $17$ with $N=1000$. Interestingly, the dynamics are super-diffusive for $V=10$, while in the nearby points, like $V=9$, it shows sub-diffusive behavior. Such behavior is not expected if the system were deep inside a stable phase. Together with the strong system-size dependence of $\gamma$ in the P4 regime (see Fig. 23), these observations suggest that the P4 phase reported in Fig. 2 is more likely an anomalous crossover regime that needs further study, which lies beyond the scope of the present work.